OpenStudy (anonymous):

Find all functions f:R→R, that satisfy f(x^3 +y^3) = x^2 f(x) + y f(y^2) for all x,y∈R

OpenStudy (experimentx):

f(x) = x ---- 1

OpenStudy (anonymous):

how to find all the functions?

OpenStudy (experimentx):

Seems somewhat like linear function ... though not exactly!!

OpenStudy (anonymous):

HI, Let\[ f(x)=x^n. Then f(x^3 +y^3) = x^2 f(x) + y f(y^2)\] is not satisfied. So f(x) can't be a polynomial. \[f(x).f(x)=x^2.y^2. NOw a = a^2 \iff a= 0 or 1\]

OpenStudy (anonymous):

@experimentX I am confused pls help!

OpenStudy (anonymous):

\[F(x+y)=F(x)+F(y)\] put \[y=0\] \[F(x)=F(x)+F(0)\] so either\[ F(x)=0 \] or \[F(0)=0 => F(x) \]has the form \[F(x)=xg(x)\] where \[g(x)=!0 \]for all \[x\] now I have to show g(x)=1 for all x. From the relations of \[F(x) and g(x)\], it can be derived \[g(x+y)=(xg(x)+yg(y))/(x+y)\] and \[g(xy)=g(x)g(y)\] put \[y=0\] in second relation \[g(x)=g(x).g(0)\] Rest @experimentX & @Davidc Do IT ! YOu CAn

OpenStudy (anonymous):

and the final answer is\[F(x) = x or F(x) = 0\]

OpenStudy (anonymous):

A medal @Davidc + @experimentX pls....

OpenStudy (anonymous):

This problem sucked me A LOT!

OpenStudy (anonymous):

the final answer is f(x)=0 ??

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

but f(x)=x is true

OpenStudy (anonymous):

i WROTE f(x)=0 OR f(x)=x

OpenStudy (experimentx):

f(x) = 0 also satisfies the given condition .... though i can't think of any other right now!!

OpenStudy (anonymous):

A medal!

OpenStudy (anonymous):

i have a idea. f(X)=k(x) , where k in R

OpenStudy (anonymous):

mention anathor question!

OpenStudy (anonymous):

Sorry Mention in other question!

OpenStudy (anonymous):

At any nonzero multiple of pi. --Elucidus

OpenStudy (experimentx):

f(x^3 +y^3) = x^2 f(x) + y f(y^2) f(x^3) = x^2 f(x) f(y^3) = yf(y^2) so x^2f(x) = xf(x^2) --> f(x^2) = xf(x) <---- not sure if i can do this!!

OpenStudy (anonymous):

Oh man Tazzz not right!

OpenStudy (anonymous):

So, f(x^3 +y^3)=f(x^3)+f(y^3)

OpenStudy (experimentx):

That a linear function ...

OpenStudy (experimentx):

f(a+b) = f(a) + (b) ... but f(a^2) = af(a) is different .. (a is a variable)

OpenStudy (anonymous):

a=x+1, f((x+1)^2)=(x+1)f(x+1).......(1) f((x+1)^2)=f(x^2+x+x+1)=f(x^2)+f(x)+f(x)+f(1)......(2)

OpenStudy (experimentx):

how can you say the function is linear??

OpenStudy (anonymous):

you say, f(x^3 +y^3) = x^2 f(x) + y f(y^2) f(x^3) = x^2 f(x) f(y^3) = yf(y^2) f(x^3 +y^3)=f(x^3)+f(y^3) f(x+y)=f(x)+f(y)

OpenStudy (experimentx):

Oh ... i got those two values putting x=0, y=0 simultaneously ... f(x^3) = x^2 f(x) f(y^3) = yf(y^2) on these grounds i can say, xf(x) = f(x^2) .... but how did you conclude that??

OpenStudy (anonymous):

f(x^3 +y^3) = x^2 f(x) + y f(y^2) =f(x^3)+f(y^3)

OpenStudy (experimentx):

\( f(x^3) = x^{3/2} f(x^{3/2})\)

OpenStudy (anonymous):

i don't understand

OpenStudy (experimentx):

f(x^2) = xf(x) not xf(x^2) = f(x^3)

OpenStudy (anonymous):

f(x^3 +y^3) = x^2 f(x) + y f(y^2) put y=0, f(x^3) = x^2 f(x)

OpenStudy (experimentx):

I don't think we could go back in reverse direction ... f(x^3+y^3) = f(x^3) = f(y^3) <---- for x=0, y=0 <--- it doesn't mean f(x^3+y^3) = f(x^3)+f(y^3) it's true only if x=0 or y=0

OpenStudy (anonymous):

f(x^3+y^3) = f(x^3) = f(y^3) is not true , but f(x^3) = x^2 f(x) for all x. f(y^3) = yf(y^2) for all y

OpenStudy (anonymous):

f(x^3 +y^3) = x^2 f(x) + y f(y^2) when we put y=0, but x can be all real numbers so it is true that f(x^3) = x^2 f(x) for all x.

OpenStudy (experimentx):

yes it must be !! so for y

OpenStudy (anonymous):

also, f(x^3 +y^3) = x^2 f(x) + y f(y^2) put x=0, but y can be all real numbers so it is true that f(y^3) = yf(y^2) for all y Thus, f satisfy f(x^3) = x^2 f(x) for all x and f(y^3) = yf(y^2) for all y

OpenStudy (experimentx):

since both x and y are domain, x^2f(x) = xf(x) can be said!!

OpenStudy (experimentx):

*x^2f(x) = xf(x^2)

OpenStudy (experimentx):

LOL ... sorry .. i posted on wrong section

OpenStudy (anonymous):

Never mind

OpenStudy (anonymous):

It is easy to show that f(x)= a x does satisfy the condition. The question is to see if the converse is true.

OpenStudy (anonymous):

Let \[ c= x^{\frac 1 3}\\ d=x^{\frac 1 3}\\ f(x+y)= f(c^3 + d^3) = f(c^3) + f(d^3)= f(x)+ f(y) \] If we know in addition that f is continuous, then \[f(x)= a \ x \]